OpenStudy (anonymous):

How do I solve this inequality? |S+4|>2

OpenStudy (aihberkhan):

Okay so this is really simple... all you have to do is subtract 4 from both sides! So you would get: \[S > -2\]

OpenStudy (aihberkhan):

Hope this helped! Have a great day! Also a medal would be much appreciated! Just click best response next to my answer. Thank You! @redskychicken

OpenStudy (anonymous):

But don't you have to flip the greater than sign to less than? @AihberKhan

TheSmartOne (thesmartone):

Your answer is wrong @AihberKhan You totally forgot to consider that there is an absolute value around S + 4

OpenStudy (anonymous):

Do you know what it is? @TheSmartOne

OpenStudy (anonymous):

I really need help with this... I got a test on this stuff tomorrow

OpenStudy (superdavesuper):

not sure if @TheSmartOne will come back n help but u can start by considering two cases: if S+4 >=0 and if S+4<0 if S+4 >=0, then |S+4| = S+4 if S+4<0, then |S+4|=-(S+4) put the above back into the original inequality then simplify.

OpenStudy (anonymous):

i got s>-2 and s<2 Would that be right? because now i have to graph it @superdavesuper

OpenStudy (superdavesuper):

if S+4 >=0, then |S+4| = S+4 > 2 subtracting 4 on left n right, u will get S>-2 if S+4<0, then |S+4|=-(S+4) > 2 multiply by -1 (n flipping the inequality) then minus 4, u will get S+4<-2 S<-6 got it? :)